Assume that A is a 3 by 3 matrix such that det(A) = −10. Let B be a matrix obtained from A using the following elementary row operations:


R3 + 2R1 → R3,

5R1 → R1,

−2R2 → R2

R2 ↔ R3.


Find the determinant of B obtained from the resulting operations, i.e., det(B)


1
Expert's answer
2021-06-02T13:58:11-0400

Answer:-



"R_3+2R_1\to R_3"

The value of a determinant is not changed by adding the elements of one column multiplied by an arbitrary number to the corresponding elements of another column.


If we multiply one row of a matrix by "t," the determinant is multiplied by "t"



"5R_1\to R_1=>t_1=\dfrac{1}{5}""-2R_2\to R_2=>t_2=-\dfrac{1}{2}"


If you exchange two rows of a matrix, you reverse the sign of its determinant from positive to negative or from negative to positive. 



"R_2\leftrightarrow R_3"

Then


"\det(B)=\dfrac{1}{5}(-\dfrac{1}{2})(-1)\det (A)=\dfrac{1}{10}(\det A)""=\dfrac{1}{10}(-10)=-1"